# Distribution Shifts at Scale: Out-of-distribution Detection in Earth Observation

### v1

@@ 6 Experimental Results
In this section, we present the ablation studies and discuss the results of the surrogate sample assignment benchmark, where we evaluate the performance of $g^{*}$. Note that ID task performance is not reported, as it remains unaffected by our method.
@@
Ablation Studies. We conduct benchmark studies on the $g_{\text{oracle}}$ classifier to explore the factors influencing TARDIS’s ability to detect test-time distribution shifts. This ability depends significantly on the layer from which internal activations are extracted and the downsampling method applied. To address this, we test different layers and compare their performance. For downsampling, we experiment with several methods, including mean and standard deviations, average pooling, max pooling, and PCA-based reduction. The max pooling-based downsampling method achieves the highest performance, likely due to its ability to retain the most salient activation patterns, which we argue is important for effective OOD detection. After selecting the optimal layer and downsampling method, we evaluate various classifiers: KNeighbors, GaussianNB, DecisionTree, ExtraTrees, LogisticRegression, SVC, RandomForestUnbalanced, RandomForest, AdaBoost, and GradientBoosting. Results indicate that Logistic Regression provides the best tradeoff between classification performance and wall time. We then compare $g^{*}$ to $g_{\text{oracle}}$ as we tune the parameters $K$ and $T$ required for surrogate label assignment. To determine $K$ and $T$, we use a Tree-structured Parzen Estimator for sampling. Specifically, we set the search boundaries for $K$ (the number of clusters) between 2 and 0.3 $\times$ M, where $M$ is the total number of samples, and for $T$ (the fraction of in-distribution samples within a cluster, below which the cluster is labeled as surrogate OOD) between 0.01 and 0.2. These values were empirically determined based on preliminary tuning for optimal performance. We then run 20 independent experiments and select the best-performing $(K,T)$ pair. Fixing $T$ based on the hyperparameter search results, we observe a recurring pattern that enables us to fix $K$ to 0.3 $\times$ M across all setups. This choice is driven by the observation that the classifier $g^{*}$ reaches performance levels close to $g_{\text{oracle}}$ when $K$ is set to 30% of the total training samples. This trend holds across both datasets, with the xBD disaster scenario (semantic shift) and EuroSAT Pasture experiment (covariate shift) showing similar behavior. In both cases, classifier performance improves as more clusters are incorporated, approaching the performance of $g_{\text{oracle}}$. We present the results of TARDIS across these controlled distribution shifts in the Supplementary Material.
@@
Surrogate Sample Assignment Benchmark. The quantitative results for $g^{*}$ and $g_{\text{oracle}}$ across 16 experimental setups are shown in Table 2. Across all setups, $g^{*}$ closely matches the performance of $g_{\text{oracle}}$, which represents the upper bound for our method. Specifically, in 13 setups for AUROC and 17 setups for FPR95, the difference between $g^{*}$ and $g_{\text{oracle}}$ is not statistically significant. In the remaining setups, while a small performance difference exists, the proposed method largely recovers the theoretical upper bound. This minimal, and often statistically insignificant, performance gap demonstrates that TARDIS can assign surrogate ID and OOD labels with high performance, even without access to test-time distribution during training.
A key takeaway from these experiments is the method’s ability to disentangle WILD samples with unknown test-time distributions into surrogate ID and OOD samples. Notably, under unknown distribution settings, it achieves performance comparable to cases with known ID and OOD labels, highlighting its strength in reliable OOD detection without prior distribution knowledge, making it suitable for real-world geospatial applications.
@@ 7 OOD Detection Goes Global: Real-World Deployment
We use the Fields of the World (FTW) pre-trained models [11] to demonstrate the capabilities of TARDIS in a large-scale deployment scenario. FTW is a geographically diverse dataset designed for agricultural field segmentation, covering 24 regions across Europe, Africa, Asia, and South America: Austria, Belgium, Brazil, Cambodia, Corsica, Croatia, Denmark, Estonia, Finland, France, Germany, India, Kenya, Latvia, Lithuania, Luxembourg, Netherlands, Portugal, Rwanda, Slovakia, Slovenia, South Africa, Spain, Sweden, and Vietnam. The dataset contains approximately 70,000 samples, each consisting of multi-date, multi-spectral Sentinel-2 satellite patches paired with three-class semantic segmentation masks (field, field boundary, and background). The task involves segmenting these classes using a pair of Sentinel-2 images — one for the planting season and one for the harvesting season — as input. Field boundary data is crucial for global agricultural monitoring, however training large scale models to segment field boundaries from satellite imagery presents significant challenges due to the geographic diversity of fields, varying crop cycles, and agro-climatic conditions, all of which which introduce substantial distribution shifts. This complexity, along with corresponding distribution shifts found in real world imagery, makes the FTW dataset ideal for testing TARDIS’s ability to diagnose model performance during inference. Additionally, the multi-date nature of the dataset is particularly suitable for evaluating segmentation models that must handle spatiotemporal variations in satellite imagery.
@@
Sampling WILD Set. We collect multispectral Sentinel-2 satellite images using the Microsoft Planetary Computer [22], which is freely accessible. The images are processed to Level-2A (bottom-of-atmosphere) and stored in cloud-optimized GeoTIFF (COG) format. We use four spectral bands: Red (B04), Green (B03), Blue (B02), and Near-Infrared (B08), each with a spatial resolution of 10 meters per pixel. We randomly query 1200 Sentinel-2 scenes from those over land and with cloud coverage $<10\%$. Then, from each scene, we generate random patches of shape 256 $\times$ 256 $\times$ 4 pixels (filtering out samples with $\geq 10\%$ zero or NaN pixels) and query both the planting and harvesting seasons for the same locations. The planting and harvesting dates are fixed based on the hemisphere: April 1 to June 30 and September 1 to November 30 for the Northern Hemisphere, and October 1 to December 31 and March 1 to May 31 for the Southern Hemisphere. The two patches are concatenated along the channel dimension, resulting in a patch of shape 256 $\times$ 256 $\times$ 8 pixels. Following the pre-trained model’s convention, the concatenated patch is upsampled by a factor of 2 to 512 $\times$ 512 $\times$ 8, processed by the model, and then downsampled back to the original dimensions. For surrogate label assignment, we fix the hyperparameters $K$ and $T$ to $0.3\times 1200$ and 0.1, respectively, as suggested by the benchmark study.
@@
The geographical distribution of ID and WILD samples is illustrated in Figure 3, where WILD samples are further classified into ID and OOD categories. A notable pattern is that samples located in arid biomes — such as the deserts in Inner Australia, the Sahara, and Patagonia — and polar regions, including Icelandic glaciers and the South Pole, are more likely to be assigned as OOD. This can be attributed to the ID samples predominantly representing mesic environments, which are areas with moderate moisture levels and managed landscapes suitable for agriculture. In contrast, arid biomes and polar regions exhibit extreme environmental conditions that differ markedly from the mesic environments of the ID samples. This ecological dissimilarity results in these biomes being classified as surrogate OOD, highlighting the model’s sensitivity to environmental context and distribution shifts.
@@
To evaluate the scalability of our method, we measure the time required to classify the internal activation of a WILD sample as surrogate ID or OOD using the classifier $g$, implemented as a logistic regressor. An input patch of shape $256\times 256$ pixels, with a spatial resolution of 10 meters per pixel, covers an area of $6.5536\,\text{km}^{2}$ and takes 0.003 seconds to process. Applied to the entirety of mainland Africa, which spans $29.77\times 10^{6}\,\text{km}^{2}$, the method could be deployed on a single CPU in approximately 15 days. In a distributed GPU setup, this process could be completed within hours (a fraction of the time required for $f$ model inference over a similar area). This is crucial as such a model would allow practitioners to see, spatially, where the field boundary model itself is performing in distribution versus not (and where the corresponding outputs are likely irrelevant). These results demonstrate the practicality of the proposed method for deployment in low-data regimes, where geospatial models are typically less robust, and highlight its potential for improving model understanding in these challenging settings.
We further evaluate the reliability of the OOD classifier $g$ on the training set by comparing the performance of the $f$ model on the FTW dataset’s test set with the probability scores from $g$. We observe a correlation between the performance of the $f$ model on the test set and low skewness in the output of $g$. This indicates that when the $f$ model performs poorly, the OOD classifier $g$ reflects this by producing unimodal OOD scores, as shown by low skewness values (see Figure 4) for countries such as Portugal, Cambodia, and Vietnam.
Figure 4: Relationship between model performance and OOD score skewness. The figure illustrates the correlation between low performance of the $f$ model on the FTW test set and low skewness in the OOD classifier $g$’s score distribution. When the $f$ model performs poorly, the OOD classifier $g$ produces more OOD scores, as reflected by lower skewness values.
@@ 8 Conclusion
We present TARDIS, which effectively assigns surrogate ID or OOD labels to samples with unknown distributions by clustering the representation space. These surrogate labels are used to train a binary classifier for OOD detection. We do not aim to optimize OOD detection performance but instead propose a method designed for real-world applications where the distribution is unknown, maintaining ID task performance is critical, and computationally expensive methods are impractical. It achieves performance close to the theoretical upper limit in assigning surrogate ID and OOD samples in 13 out of 17 experiments. We demonstrate TARDIS can be deployed at scale, with the potential to enhance the robustness and trustworthiness of geospatial models as well as provide insights into distribution shifts in global applications.
@@ 9.1 EuroSAT
We simplify the damage assessment task into binary segmentation by reassigning damage levels: background (0) and levels 1-2 are grouped, while levels 3-4 form a high-damage class. This minimizes concept drift and ensures a balanced evaluation. We train a U-Net model with a ResNet50 backbone, pre-trained on ImageNet and configured for 3 input channels. Training uses a batch size of 32, a learning rate of 0.001, and runs for up to 50 epochs with early stopping after 5 epochs of no improvement. We reserve 10% of the data for validation and normalize the input images by dividing pixel values by 255.
@@ 10.1 Which Layer?
Figure 6: Examples from the dataset, with one sample from each class. These images highlight the spatial and semantic distinctions across classes.
Figure 7: The input samples are from the WILD set, where the distribution is unknown. The figure shows Sentinel-2 images at two different times (planting season and harvesting season — Window A and Window B). When these windows are fed together into $f$, the model outputs both the segmentation prediction and the OOD classifier $g$’s prediction score.
@@ 10.2 Which Downsampling Method?
    PCA-based reduction (PCA): Applies Principal Component Analysis (PCA) to reshape the activation map into a vector and projects it into a lower-dimensional space with 10 components.
@@ 10.4 Surrogate Label Assignment: Hyperparameter Search for k and T
The underlying assumption is that samples with similar distributions lie closer in the activation space than those from dissimilar distributions. Effectively segmenting the activation space is critical, as the distributions of WILD samples are unknown during deployment. Segmenting nearby activations and correctly assigning clusters to surrogate ID or OOD makes the choice of k and T crucial for accurate OOD detection.
@@
The primary goal is to understand how to choose k and T, and whether there are patterns we can extrapolate to real-life deployment. For this, we calculate the ratio of k to the total number of training samples and evaluate its effect on OOD detection performance (Accuracy, FPR95, and AUROC). We plot these metrics against the ratio of k/total training samples, increasing k until the ratio reaches 1. Theoretically, OOD detection improves with more clusters as this enables finer-grained segmentation of the activation space, reducing the risk of including anomalies in ID clusters.
To establish a theoretical maximum (oracle performance), we also evaluate OOD classification with known ID and OOD labels, bypassing the need for clustering. This oracle performance is represented by horizontal dashed lines in Figures 8 and 9 (upper plots). The results for two representative experiments—one from EuroSAT and one from xBD—since all experiments show similar trends. We observe that the performance approaches the oracle boundaries when k is approximately 0.3 times the total number of training samples. While performance improves as k increases, a trade-off is required between performance and walltime as well as computational complexity. Based on this trade-off, we set k to 0.3 for all experiments, including the large-scale deployment on FTW. Furthermore, we observe that our method is not highly sensitive to T. As a result, we fix T to 0.1 for all experiments, which is the value used in this initial investigation. We use the Optuna library to implement a Bayesian-based search algorithm. The composite objective function, which we minimize to determine the optimal number of clusters and ID fraction threshold, is detailed in Section 4.
@@
We set k and T as described and use t-SNE in the lower plots of Figures 8 and 9 to reduce the dimensionality of the activation spaces to 2D for visualization. When ID and OOD labels are known, the t-SNE plots show that only a small fraction of labels changes from the original labels. This demonstrates the effectiveness of the surrogate label assignment process described above.
11 Additional Experimental Results
In Figure 10, we show the predictions of the DL model $f$ and the OOD classifier $g$, along with the ground truth class and distribution annotations for the EuroSAT experiment, where Forest serves as the OOD class. The model $f$ trains on 9 classes (excluding Forest) and tests on Forest. The first row shows correct predictions by $f$, while the second row shows incorrect predictions. Even when $f$ makes misclassifications, $g$ accurately quantifies the distribution shifts in most cases. The performance of $f$ on the test set is not directly measurable since the test uses a single unseen class. We report the performance of $g$ as: Accuracy = 0.9325, ROC AUC = 0.9886, and FPR95 = 0.0619.
For xBD, we present results where $f$ trains on Hurricane Matthew (ID, Figure 11) and tests on Mexico Earthquake (OOD, Figure 12). Comparing the input images and masks between ID and OOD shows that even when $f$ performs suboptimally, $g$ effectively quantifies the distribution shifts. The performance of $f$ on the test set is: Multiclass Accuracy = 0.7690 and Multiclass Jaccard Index = 0.6248. We attribute $f$’s suboptimal prediction performance to the significant distribution shift between the training (Hurricane Matthew) and testing (Mexico Earthquake) datasets. The performance of $g$ is: Accuracy = 0.9806, ROC AUC = 0.9986, and FPR95 = 0.0.

### v2

@@ 6 Experimental Results
In this section, we present ablation studies and experimental results. Note that ID task performance is not reported, as it remains unaffected by our method.
@@
Ablation Studies. We conduct benchmark studies on the $g_{\text{oracle}}$ classifier to explore the factors influencing TARDIS’s ability to detect test-time distribution shifts. This ability depends significantly on the layer from which internal activations are extracted and the downsampling method applied. To address this, we test different layers and compare their performance (Table 4 and Table 5). For downsampling, we experiment with several methods, including mean and standard deviations, average pooling, max pooling, and PCA-based reduction. The max pooling-based downsampling method achieves the highest performance, likely due to its ability to retain the most salient activation patterns, which we argue is important for effective OOD detection (Table 6). We then evaluate various classifiers: KNeighbors, GaussianNB, DecisionTree, ExtraTrees, LogisticRegression, SVC, RandomForestUnbalanced, RandomForest, AdaBoost, and GradientBoosting. Results indicate that Logistic Regression provides the best tradeoff between classification performance and wall time (Table 7). We then compare $g^{*}$ to $g_{\text{oracle}}$ as we tune the parameters $k$ and $T$ required for surrogate label assignment. To determine $k$ and $T$, we use a Tree-structured Parzen Estimator for sampling. Specifically, we set the search boundaries for $k$ (the number of clusters) between 2 and 0.3 $\times$ $M$, where $M$ is the total number of samples, and for $T$ (the fraction of in-distribution samples within a cluster, below which the cluster is labeled as surrogate OOD) between 0.01 and 0.2. These values were empirically determined based on preliminary tuning for optimal performance. We then run 20 independent experiments and select the best-performing $(k,T)$ pair. Fixing $T$ based on the hyperparameter search results, we observe a recurring pattern that enables us to fix $k$ to 0.3 $\times$ M across all setups. This choice is based on the observation that the classifier $g^{*}$ reaches performance levels close to $g_{\text{oracle}}$ when $k$ is set to 30% of the training samples. This trend holds for both datasets, with the xBD disaster scenario (semantic shift) and the EuroSAT Pasture experiment (covariate shift) showing similar behavior. In both cases, performance improves as more clusters are added, approaching that of $g_{\text{oracle}}$. In Figures 8 and 9, we show how we determine the optimal $(k,T)$ pair that achieves high surrogate label assignment performance across all setups. The same pair generalizes to real-world deployment, removing the need for hyperparameter tuning.
@@
Surrogate Sample Assignment Benchmark. Table 2 shows results for $g^{*}$ and $g_{\text{oracle}}$ across 17 setups. Overall, $g^{*}$ closely approaches $g_{\text{oracle}}$, which serves as the upper bound. In 13 setups for AUROC and all 17 for FPR95, the performance gap is not statistically significant. In the few remaining cases, the gap is small. These results show that TARDIS can effectively assign surrogate labels without access to test-time labels.
ID/OOD Balance. As discussed in Section 4, splitting ID and WILD data into validation and clustering sets affects the ID/OOD ratio throughout the pipeline. In Table 2, WILD consists entirely of OOD samples, but in real deployments, it may contain both ID and OOD. While data is shuffled, monitoring this ratio is key to understanding sensitivity. For example, in the Pasture semantic shift setup, the clustering and classifier stages see 17% and 15% OOD, respectively. In Nepal Flood – Midwest Flood, the ratios are 46% and 35%. This variation shows that TARDIS performs robustly across different ID/OOD balances, supporting the generalizability of our method.
Figure 3: Geographical distribution of ID and WILD sets, containing 500 and 1200 samples, respectively. The ID set is sampled from the FTW dataset training set, while the WILD set is randomly sampled from the Microsoft Planetary Computer. Each Sentinel-2 patch, provided in two different time frames (planting and harvesting). The model $f$ takes both Sentinel-2 images from different seasons as input, and the predictions are shown on the right. The OOD probability values are thresholded at 0.5.
Table 3: Performance comparison of post-hoc score-based and activation-based OOD detection methods across 10 semantic shift and 7 covariate shift experiments on EuroSAT and xBD. We report the mean ± standard deviation across all 10 or 7 experiments for each method. TARDIS achieves comparable performance while requiring no hyperparameter tuning and preserving main task performance.
[…]
Semantic Shift Mahalanobis [19] 0.85 $\pm$ 0.01 0.19 $\pm$ 0.05
MSP [13] 0.90 $\pm$ 0.03 0.23 $\pm$ 0.07
Energy [21] 0.89 $\pm$ 0.02 0.20 $\pm$ 0.04
ReAct [30] 0.92 $\pm$ 0.04 0.13 $\pm$ 0.01
NAP [25] 0.93 $\pm$ 0.03 0.21 $\pm$ 0.05
TARDIS (Ours) 0.95 $\pm$ 0.02 0.16 $\pm$ 0.03
Covariate Shift Mahalanobis [19] 0.88 $\pm$ 0.03 0.22 $\pm$ 0.06
MSP [13] 0.92 $\pm$ 0.04 0.25 $\pm$ 0.08
Energy [21] 0.91 $\pm$ 0.03 0.22 $\pm$ 0.05
ReAct [30] 0.94 $\pm$ 0.05 0.15 $\pm$ 0.02
NAP [25] 0.95 $\pm$ 0.02 0.23 $\pm$ 0.06
TARDIS (Ours) 0.98 $\pm$ 0.01 0.11 $\pm$ 0.04
Comparison with Existing OOD Detection Methods. We benchmark existing post-hoc activation-based and score-based methods across all semantic and covariate shift experiments, reporting mean and standard deviation (Table 3). Gradient-based methods involving backpropagation are excluded due to their high computational cost and increased wall time. Score-based methods (Mahalanobis, MSP, Energy) preserve ID performance as they require no model changes, but rely on potentially overconfident predictions and are less sensitive to near-domain shifts. Activation-based methods like ReAct and NAP modify internal activations—via clipping or purification—boosting OOD detection but at the cost of reduced ID performance, limiting their practicality.
Why Clustering Alone is Insufficient. Our method includes a distance-based clustering step followed by an additional classification step (Step 3 and 4 in Figure 2). Removing the classification step results in a drop in ROC AUC for the River experiment from 97% to 79%—a trend consistently observed across all experiments. This demonstrates that clustering alone (i.e., distance-based assessment) is insufficient for effective distribution detection, justifying the need for Step 4 in Figure 2.
@@ 7 OOD Detection Goes Global: Real-World Deployment
We use the Fields of the World (FTW)  [14] training set and pre-trained models to demonstrate the capabilities of TARDIS in a large-scale deployment scenario. FTW is a geographically diverse dataset designed for agricultural field segmentation, covering 24 regions across Europe, Africa, Asia, and South America. The dataset contains approximately 70,000 samples, each consisting of multi-date, multi-spectral Sentinel-2 satellite patches paired with three-class semantic segmentation masks (field, field boundary, and background). The task involves segmenting these classes using a pair of Sentinel-2 images — one for the planting season and one for the harvesting season — as input. Field boundary data is crucial for global agricultural monitoring, however training large scale models to segment field boundaries from satellite imagery presents significant challenges due to the geographic diversity of fields, varying crop cycles, and agro-climatic conditions, all of which introduce substantial distribution shifts. This complexity, along with corresponding distribution shifts found in real world imagery, makes the FTW dataset ideal for testing TARDIS’s ability to diagnose model performance during inference. Additionally, the multi-date nature of the dataset is particularly suitable for evaluating models that must handle spatiotemporal variations in satellite imagery.
@@
Sampling WILD Set. We collect multispectral Sentinel-2 L2A satellite images using the Microsoft Planetary Computer [28]. The images are processed to Level-2A (bottom-of-atmosphere) and stored in cloud-optimized GeoTIFF (COG) format. We use four spectral bands: Red (B04), Green (B03), Blue (B02), and Near-Infrared (B08), each with a spatial resolution of 10 meters per pixel. We randomly query 1200 Sentinel-2 scenes from those over land and with cloud coverage $<10\%$. Then, from each scene, we generate random patches of shape  256 $\times$ 256 $\times$ 4 pixels (filtering out samples with $\geq 10\%$ zero or NaN pixels) and query both the planting and harvesting seasons for the same locations. The planting and harvesting dates are fixed based on the hemisphere: April 1 to June 30 and September 1 to November 30 for the Northern Hemisphere, and October 1 to December 31 and March 1 to May 31 for the Southern Hemisphere. The two patches are concatenated along the channel dimension, resulting in a patch of shape 256 $\times$ 256 $\times$ 8 pixels. Following the pre-trained model’s convention, the concatenated patch is upsampled to 512 $\times$ 512 $\times$ 8, processed by the model, and then downsampled back to the original dimensions. For surrogate label assignment, we fix the hyperparameters $k$ and $T$ to $0.3\times 1200$ and 0.1, respectively, as suggested by the benchmark study.
@@
The geographical distribution of ID and WILD samples is shown in Figure 3, with WILD samples further classified into surrogate ID and OOD. A clear pattern emerges: samples from arid biomes—such as the deserts of Inner Australia, the Sahara, and Patagonia—and polar regions, including Icelandic glaciers and the South Pole, are more frequently assigned as OOD. This is likely because the ID samples predominantly represent mesic environments: moderately moist, managed landscapes typical of agricultural areas. In contrast, arid and polar regions exhibit extreme environmental conditions that differ significantly from those in the ID set. This ecological dissimilarity likely drives their classification as surrogate OOD, reflecting the model’s sensitivity to environmental context and distribution shifts.
@@
To evaluate scalability, we measure the time required to classify a WILD sample’s internal activation as surrogate ID or OOD using the logistic regression classifier $g$. A $256\times 256$ input patch (10 m resolution, covering $6.5536\,\text{km}^{2}$) takes 0.003 seconds to process. Applied to mainland Africa ($29.77\times 10^{6}\,\text{km}^{2}$), the method runs on a single GPU in under 4 hours—a fraction of the time needed for full $f$-model inference. This enables spatial diagnostics, identifying where the field boundary model is likely in- or out-of-distribution and where outputs may be unreliable. Such rapid screening is especially useful in low-data regions, where models are typically less robust and OOD detection can guide targeted interventions or data collection.
We assess the reliability of the OOD classifier $g$ by comparing its output skewness on the training set to the performance of the model $f$ on the FTW test set. High skewness indicates confident (mostly ID or OOD) predictions, while low skewness suggests uncertainty or misclassification. Plotting skewness against $f$’s test performance reveals a trend: low skewness aligns with poor model performance (Figure 4), as seen in countries like Portugal, Cambodia, and Vietnam. This suggests that low skewness may signal the presence of OOD samples. Thus, TARDIS can also serve as a sanity check for detecting distribution shifts in test splits.
Limitations and Future Work. Our method assumes access to a non-trivial number of WILD samples and requires periodic updates to the clustering used for surrogate label assignment. Future work could explore density-based clustering methods instead of the current distance-based approach, as they may better handle variable sample densities and outliers, further improving surrogate label assignment.
Figure 4: Correlation between model performance and OOD score skewness. Low performance of the $f$ model on the FTW test set aligns with low skewness in the OOD classifier $g$’s score distribution, suggesting the presence of OOD samples in the test set.
@@ 8 Conclusion
We present TARDIS, a distribution shift detector that assigns surrogate ID and OOD labels to samples from unknown distributions by leveraging clustering in the feature space. These surrogate labels are assigned based on known ID samples and used to train a binary classifier that outputs a distribution score during inference. Rather than optimizing for OOD detection performance, TARDIS targets real-world scenarios where the distribution is unknown, ID task performance must be preserved, and computational overhead is a concern. Across 13 of 17 experiments mimicking real-world semantic and covariate distribution shifts, our method achieves near-upper-bound accuracy for surrogate label assignment while matching the performance of top post-hoc OOD detection methods. We demonstrate how TARDIS scales in a real-world application, providing interpretable insights into model behavior under distribution shifts and helping identify potential biases and limitations in the dataset. This contributes to the robustness and trustworthiness of models, making it particularly valuable in safety-critical, time-sensitive, and low-data settings.
@@ 9.1 EuroSAT
We simplify the damage assessment task into binary segmentation by reassigning damage levels: background (0) and levels 1-2 are grouped, while levels 3-4 form a high-damage class. This minimizes concept drift and ensures a fair evaluation of distribution shifts. We train a U-Net model with a ResNet50 backbone, pre-trained on ImageNet and configured for 3 input channels. Training uses a batch size of 32, a learning rate of 0.001, and runs for up to 50 epochs with early stopping after 5 epochs of no improvement. We reserve 10% of the data for validation and normalize the input images by dividing pixel values by 255.
@@ 10.1 Which Layer?
Figure 6: Examples from the EuroSAT dataset, with one sample from each class. These images highlight the spatial and semantic distinctions across classes.
Figure 7: Deploying TARDIS over FTW dataset: The input samples are from the collected WILD set, where the distribution is unknown. The figure shows Sentinel-2 images at two different times (planting season and harvesting season — Window A and Window B). When these windows are fed together into $f$, the model outputs both the segmentation prediction and the OOD classifier $g$’s prediction score.
@@ 10.2 Which Downsampling Method?
    PCA-based reduction (PCA): Applies Principal Component Analysis to reshape the activation map into a vector and projects it into a lower-dimensional space with 10 components.
@@ 10.4 Surrogate Label Assignment: Hyperparameter Search for k and T
The underlying assumption is that samples with similar distributions lie closer in the activation space than those from dissimilar distributions. Effectively clustering the activation space is critical, as the distributions of WILD samples are unknown during deployment, and it depends on the optimal choice of $k$ and $T$.
@@
The primary goal is to understand how to choose k and T, and whether there are patterns we can extrapolate to real-life deployment. For this, we first assign surrogate labels and calculate the ratio of k to the total number of training samples and evaluate its effect on OOD detection performance (Accuracy, FPR95, and AUROC). We plot these metrics against the ratio of k/total training samples, increasing k until the ratio reaches 1. Theoretically, OOD detection improves with more clusters as this enables finer-grained clustering of the activation space, reducing the risk of including anomalies in ID clusters.
To establish a theoretical maximum (upper-bound performance), we also evaluate OOD classification with known ID and OOD labels, bypassing the need for clustering. This oracle performance is represented by horizontal dashed lines in Figure 8 and Figure 9 (upper plots). The results for two representative experiments—one from EuroSAT and one from xBD—since all experiments show similar trends. We observe that the performance approaches the oracle boundaries when k is approximately 0.3 times the total number of training samples. While performance improves as k increases, a trade-off is required between performance and walltime as well as computational complexity. Based on this trade-off, we set k to 0.3 for all experiments, including the large-scale deployment on FTW. Furthermore, we observe that our method is not highly sensitive to T. As a result, we fix T to 0.1 for all experiments, which is the value used in this initial investigation. We use the Optuna library to implement a Bayesian-based search algorithm. The composite objective function, which we minimize to determine the optimal number of clusters and ID fraction threshold, is detailed in Section 4.
@@
We set k and T as described and use t-SNE in the lower plots of Figure 8 and Figure 9 to reduce the dimensionality of the activation spaces to 2D for visualization. When ID and OOD labels are known, the t-SNE plots show that only a small fraction of labels changes from the original labels. This demonstrates the effectiveness of the surrogate label assignment process described above.
11 Further Experimental Results
In Figure 10, we show the predictions of the DL model $f$ and the OOD classifier $g$, along with the ground truth class and distribution annotations for the EuroSAT experiment, where Forest serves as the OOD class. The model $f$ trains on 9 classes (excluding Forest) and tests on Forest. The first row shows correct predictions by $f$, while the second row shows incorrect predictions. Even when $f$ makes misclassifications, $g$ accurately quantifies the distribution shifts in most cases. The performance of $f$ on the test set is not directly measurable since the test uses a single unseen class. We report the performance of $g$ as: Accuracy: $93.25\%$, ROC AUC: $98.86\%$, FPR95: $6.19\%$.
For xBD, we present results where $f$ is trained on Hurricane Matthew (ID, Figure 11) and tested on Mexico Earthquake (OOD, Figure 12). Comparing the input images and masks between ID and OOD reveals that even when $f$ performs suboptimally, $g$ effectively quantifies the distribution shifts. The performance of $f$ on the test set is as follows: Multi-class accuracy: $76.90\%$, Multi-class Jaccard index: $62.48\%$. We attribute $f$’s suboptimal prediction performance to the significant distribution shift between the training (Hurricane Matthew) and testing (Mexico Earthquake) datasets, and also to the fact that we reformulate the main task of damage classification to building detection (as described in Section 9.2). The performance of $g$ is: Accuracy: $98.06\%$, ROC AUC: $99.86\%$, FPR95: $0.00\%$.